# Layered Insights: Generalizable Analysis of Authorial Style by Leveraging All Transformer Layers

### v1

@@ Abstract
We propose a new approach for the authorship attribution task that leverages the various linguistic representations learned at different layers of pre-trained transformer-based models. We evaluate our approach on three datasets, comparing it to a state-of-the-art baseline in in-domain and out-of-domain scenarios. We found that utilizing various transformer layers improves the robustness of authorship attribution models when tested on out-of-domain data, resulting in new state-of-the-art results. Our analysis gives further insights into how our model’s different layers get specialized in representing certain stylistic features that benefit the model when tested out of the domain.
**footnotetext: These authors contributed equally to this work
@@ 1 Introduction
Authorship attribution (AA) is the task of analyzing the writing style of pairs of texts in order to predict whether they are written by the same author. Approaches developed for this task either focus on explicitly modeling linguistic attributes of texts Koppel and Schler (2004) or rely on pre-trained transformer-based models to learn relevant features to the task from big training corpora Rivera-Soto et al. (2021); Wegmann et al. (2022). The latter approach achieves state-of-the-art results on the AA task. However, it focuses on modeling texts by only learning from the final output layer of the transformer, ignoring representations learned at other layers. A large body of research has been investigating the inner workings of pre-trained transformer-based models Rogers et al. (2020), showing that different layers are specialized in modeling different linguistic phenomena of texts, with lower layers mainly modeling lexical features while higher ones model sentence structure Rogers et al. (2021).
@@
Figure 1: Our approach is a transformer-based model where the initial embeddings (Init. Embed.) and the output of each layer passes through a projection layer, obtaining N+1 document representations that are jointly optimized on the authorship attribution task using contrastive learning
@@
In this paper, we hypothesize that leveraging the representations learned at different layers of a pre-trained transformer can lead to more effectiveness on the authorship attribution task, especially since the AA task requires modeling similarities between different linguistic aspects of the texts. For example, the writing style of an author can be realized by their use of similar vocabulary (lexical), sentence structure (syntax), or even the overall discourse they follow in their writing. To study this hypothesis, we develop a new approach to the task that learns to capture these different stylistic aspects by leveraging all the transformer layers. In particular, as highlighted in Figure 1, starting from a pre-trained transformer-based model with N layers, our approach projects the initial token embeddings as wel as each of the N layers’ embedding into a new embedding space, where we apply a contrastive loss that refines this space to better capture the similarity between texts with respect to the original representation. The model is then trained jointly by averaging the contrastive losses at all the N+1 projections. During prediction, given two texts, the analysis of whether they belong to the same author can be decomposed into comparing their similarities at these different learned projections, leveraging more robust predictive power by aggregating these scores.
@@
To evaluate our hypothesis, we follow the same evaluation setup of Rivera-Soto et al. (2021). We train and evaluate our approach against state-of-the-art models on three datasets: Reddit comments, Amazon reviews, and Fanfiction short stories. Our experiments demonstrate that, for the out-of-domain scenario, our approach performs best in all settings, with the highest increase of around 50% in the mean reciprocal rank (MRR) for our model trained on the Fanfiction dataset and evaluated on Reddit. Results, however, in the in-domain scenario are mixed. These insights indicate strong generalization capabilities of our model compared to baselines, which we argue is more relevant in practical scenarios of the AA task than in-domain. This also provides evidence of the importance of leveraging the various layers of the base transformer when working on the AA task. Furthermore, we perform experiments to analyze the role each layer plays in the overall performance of our model. We found that certain layers became more relevant when our model tested in the out-of-domain setting, leading to high performance.
@@ 2 Related Work
Author Attribution. Traditional methods rely on manually designed features, such as syntactic patterns and function word frequencies, to capture distinctive writing styles Koppel and Schler (2004). With the rise of deep learning, transformer-based models have set new records by leveraging their ability to capture nuanced stylistic patterns in text. For instance, Rivera-Soto et al. (2021) employed a contrastive learning approach to map texts written by the same author into similar regions of an embedding space. Similarly, Wegmann et al. (2022) introduced methods to ensure that embeddings focus on style rather than content, allowing for more reliable attribution. Despite their success, these approaches rely only on the final layer of the transformer model as the source of text representation. In contrast, our approach leverages different text representations from all the transformer’s layers.
@@ 3 Approach
Given mounting evidence from the literature on the various linguistic capabilities of the different layers of transformer-based models, we propose a multi-layer contrastive learning approach that explicitly models authorship attribution at varying levels of linguistic granularity. Instead of relying on a single representation derived from the final layer, our approach extracts embeddings from all the N layers of the transformer to leverage its linguistic representations for the authorship attribution task.
As highlighted in Figure 1, our model is RoBERTa-based Liu (2019) with additional N+1 Feed Forward layers that project each of the hidden states of RoBERTa into a new embedding space, including the initial token embeddings. We then apply a contrastive loss function on each of these N+1 embedding spaces to pull documents written by the same author together while keeping other documents further apart in the corresponding space. At each training step, the average loss of all the N contrastive losses is computed and back-propagated throughout the model. We hypothesize that this architecture ensures that stylistic features across multiple linguistic levels contribute to the final representation, allowing the model to learn distinct and complementary signals from different layers. Rather than collapsing stylistic information into a single embedding, we maintain separate representations for each layer. In Section 5.1, we present an analysis confirming how the various learned layers of our trained model capture different linguistic representations from texts. During inference, given two texts, we compute their similarity at each learned layer projection. These layer-wise similarity scores are aggregated to produce a final similarity measure that determines the likelihood of these two texts being written by the same author.
4 Evaluation
We first describe the experiment setup to evaluate our model’s performance on authorship attribution, then present our results.
@@ 4 Evaluation
Table 1: Evaluation results comparing our approach (MultiLUAR) against other baselines on the three datasets for in-domain (highlighted grey) and out-of-domain settings. The highest values are marked bold.
4.1 Experiment setup
In our evaluation, we closely follow the experiment setup of Rivera-Soto et al. (2021) to reproduce their results and compare our approach against theirs.
@@ 4.1 Experiment setup
We evaluate the effectiveness of our model on the Reddit, Amazon and Fanfiction datasets used in Rivera-Soto et al. (2021). These datasets represent distinct domains with different linguistic styles, levels of formality, and authorial consistency. Each dataset is partitioned into training and evaluation, ensuring that evaluation texts are temporally disjoint from training data where timestamps are available Andrews and Bishop (2019). For Amazon, we use 135K authors with at least 100 reviews, training on 100K, and evaluating on 35K Ni et al. (2019). The Fanfiction dataset Bevendorff et al. (2020) consists of 278,169 stories from 41K authors, with evaluation restricted to 16,456 authors, each contributing exactly two stories Bischoff et al. (2020). The Reddit dataset includes 1 million authors, with training data from Khan et al. (2021) and evaluation using a disjoint, temporally future set Baumgartner et al. (2020). Further information regarding the datasets can be found in the appendix.
@@
We evaluate all models using Recall-at-8 (R@8) and Mean Reciprocal Rank (MRR). Recall-at-8 measures the probability that the correct author appears among the top 8 ranked candidates, while MRR evaluates ranking quality based on the position of the first correct author match. Higher values in both metrics indicate stronger performance Voorhees and Tice (2000).
Baselines.
@@
We compare our approach against the following baseline models trained on the same datasets. The primary baseline is the model proposed by Rivera-Soto et al. (2021), called LUAR. This baseline is a contrastive learning-based authorship attribution model with a RoBERTa model (paraphrase-distilroberta checkpoint [^1] https://huggingface.co/distilbert/distilroberta-base) as its base transformer. Given the two text collections, the model samples excerpts of 32 tokens and passes them through a RoBERTa transformer, obtaining only the final layer representation of each excerpt. An attention mechanism and max pooling layer are applied to aggregate the extracted representations into a single embedding representing each text collection. A contrastive loss is then applied to these representations to make them closer for positive pairs (written by the same author) and further apart for negative pairs. Additionally, we evaluate against the same baselines presented in Rivera-Soto et al. (2021). The first is a convolutional model Andrews and Bishop (2019), which encodes text representations using subword convolutional networks rather than transformers. The second baseline is a TF-IDF baseline, which represents documents as bag-of-words vectors weighted by term frequency in the text and normalized by inverse document frequency. We follow the training process of Rivera-Soto et al. (2021) to train a reproduced version of their LUAR model while we report the performance of the other two baselines as mentioned by the authors.
Our Model.
Our model extends the original LUAR architecture mentioned above by utilizing multiple transformer layers instead of a single final-layer embedding. Each layer’s hidden states are projected into separate embedding spaces and optimized using contrastive loss, capturing stylistic features across different linguistic levels. As a result, our model has seven projection layers: Layer 0 is the projection of the initial token embeddings of paraphrase-distilroberta transformer [^2] We also experimented with excluding the initial token embeddings, but the performance is worse. See Table 3 in the appendix, and layers 1 to 6 are the projection of its six hidden layers. We refer to our model as MultiLUAR. Following the LUAR training paradigm, we train a model for each of the three datasets for 20 epochs, respectively, with validation after every five epochs. We use a batch size of 84 and sample from every author 16 text excerpts of 32 tokens each. We evaluate two scenarios: in-domain (same dataset for training and evaluation) and out-of-domain (trained on one dataset, evaluated on other out-of-domain datasets). This setup allows us to assess the model’s generalization ability across different writing styles and domains, a scenario that has proven to be harder to achieve strong results in. Further details on the training process can be found in the appendix.
@@
Training Dataset Evaluation Dataset Model Reddit $\cup$ Fanfic Reddit $\cup$ Amazon Amazon $\cup$ Fanfic R@8 MRR R@8 MRR R@8 MRR Reddit Ours 64.13 49.46 63.96 49.44 29.95 20.62 LUAR 56.35 41.20 60.58 45.46 20.40 13.11 Amazon Ours 69.43 57.04 89.00 78.38 85.99 74.23 LUAR 53.51 39.63 84.84 72.09 79.60 64.83 Fanfic Ours 54.16 44.56 43.24 33.54 51.87 42.63 LUAR 57.51 46.32 40.69 30.49 51.84 41.79
@@
Table 2: Evaluation Results after training on multiple domain datasets
4.2 Results
Similar to Rivera-Soto et al. (2021), we examine the performance of our model when trained on single as well as multiple domains. We present our findings in the following sections.
4.2.1 Single-Domain Training
We first evaluate our MultiLUAR model in the single-domain setting, where models are trained on a single dataset and tested on the test split of all three datasets to assess their generalizability. Table 1 shows the recall at 8 (R@8) and mean reciprocal rank (MRR) scores of all models trained and evaluated on the different datasets. In general, LUAR and our MultiLUAR model outperform the other two baselines in all scenarios. However, for in-domain evaluation, we observe that our approach performs better than LUAR only on the Amazon dataset, while it performs slightly worse in the case of Reddit and Fanfiction datasets. This suggests that for in-domain evaluation, final-layer representations contain sufficient information for authorship attribution, and incorporating multiple layers might lead to a slight degradation in the performance. We argue that the improvement on the Amazon dataset is due to the structured writing style allowing the different layers to capture more nuanced linguistic representations from texts. In contrast, informal and discussion-based datasets, such as Reddit, exhibit less improvement, likely due to the topic-driven nature of the writing, which can overshadow stylistic signals.
For out-of-domain evaluation, our MultiLUAR approach demonstrates substantial improvements over LUAR baseline, significantly enhancing cross-domain generalization and establishing new state-of-the-art results. MultiLUAR trained on homogeneous datasets, such as Amazon, show remarkable adaptability to free-form writing (An MRR of 18.03 for MultiLUAR compared to 14.65 for LUAR on Reddit dataset), while those trained on Reddit generalize exceptionally well to both Amazon and Fanfiction. Compared to LUAR, our model substantially improves transferability, highlighting the importance of leveraging different linguistic granularity levels rather than relying solely on final-layer representations. Most notably, in the Fanfiction transfer setting, where LUAR struggles, our approach achieves 50% R@8 improvements, effectively capturing deep stylistic signals that enhance authorship attribution even in highly creative writing domains. These results confirm that our approach enhances robustness, making our approach more effective for authorship attribution across diverse linguistic contexts.
4.2.2 Multi-domain Training
Additionally, we replicate the same experiment in Rivera-Soto et al. (2021) where we train LUAR and our MultiLUAR model on two domains and evaluate them on the third domain to examine whether exposure to multiple domains enhances MultiLUAR’s robustness vs LUAR. To this end, we train our approach on dataset combinations and evaluate them on individual domains. We train models on domain pairs using mini-batches of 256 that are randomly selected from document collections, ensuring equal representation from both domains. Since authors are disjoint across domains, shared authorship occurs only within the same domain.
@@ 4.2.2 Multi-domain Training
Figure 2: The linguistic capability of each of the 6 layers of paraphrase-distilroberta model (right side) and MultiLUAR trained on Reddit dataset (left side), as probed using Holmes explorer.

### v2

# Layered Insights: Generalizable Analysis of Human Authorial Style by Leveraging All Transformer Layers

@@ Abstract
We propose a new approach for the authorship attribution task that leverages the various linguistic representations learned at different layers of pre-trained transformer-based models. We evaluate our approach on two popular authorship attribution models and three evaluation datasets, in in-domain and out-of-domain scenarios. We found that utilizing various transformer layers improves the robustness of authorship attribution models when tested on out-of-domain data, resulting in new state-of-the-art results. Our analysis gives further insights into how our model’s different layers get specialized in representing certain stylistic features that benefit the model when tested out of the domain.
@@ 1 Introduction
Identifying the author of a given text is important for various applications, such as forensic investigation and intellectual property. Authorship attribution (AA) is the task of analyzing the writing style of pairs of texts in order to predict whether they are written by the same author. Approaches developed for this task either focus on explicitly modeling linguistic attributes of texts Koppel and Schler (2004) or rely on pre-trained transformer-based models to learn relevant features to the task from large training corpora Rivera-Soto et al. (2021); Wegmann et al. (2022). The latter approach achieves state-of-the-art results on the AA task. However, it focuses on modeling texts by only learning from the final output layer of the transformer, ignoring representations learned at other layers. A large body of research has been investigating the inner workings of pre-trained transformer-based models Rogers et al. (2020), showing that different layers are specialized in modeling different linguistic phenomena of texts, with lower layers mainly modeling lexical features while higher ones model sentence structure Rogers et al. (2021).
@@
Figure 1: Our approach is a transformer-based model where the initial embeddings (Init. Embed.) and the output of each layer pass through a projection layer, obtaining N+1 document representations that are jointly optimized on the authorship attribution task using contrastive learning
@@
In this paper, we hypothesize that leveraging the representations learned at different layers of a pre-trained transformer can lead to more effective models of AA, especially since the AA task requires modeling similarities between different linguistic aspects of the texts. For example, an author’s writing style can be realized by their use of similar vocabulary (lexical), sentence structure (syntax), or even the overall discourse they follow in their writing. To study this hypothesis, we develop a new approach to the task that we call LIGHT, providing layered insights for the generalizable analysis of human authorial style by leveraging all transformer layers. In particular, as highlighted in Figure 1, starting from a pre-trained transformer-based model with N layers, our approach projects the initial token embeddings as well as each of the N layers’ embedding into a new embedding space, where we apply a contrastive loss that refines this space to better capture the similarity between texts with respect to the original representation. The model is then trained jointly by averaging the contrastive losses at all the N+1 projections. During prediction, given two texts, the analysis of whether they belong to the same author can be decomposed into comparing their similarities at these different learned projections, leveraging more robust predictive power, and allowing a fine-grained analysis of text similarities, a step towards a more interpretable approach to authorship attribution.
@@
To evaluate our hypothesis, we follow the same evaluation setup of Rivera-Soto et al. (2021) that considers three datasets for evaluation: Reddit comments, Amazon reviews, and Fanfiction short stories. We apply our approach to two popular AA models: LUAR Rivera-Soto et al. (2021) and Wegmann Wegmann et al. (2022), considering two base transformers for each: RoBERTa Liu (2019) and ModernBERT Warner et al. (2024). Our experiments demonstrate that, for the out-of-domain scenario, applying our approach on two AA models with two base transformers leads to improvement in 15 out of 16 scenarios, with the highest improvement of around 50% in the mean reciprocal rank (MRR) for LUAR when trained on the Fanfiction dataset and evaluated on Reddit. Results, however, in the in-domain scenario are mixed. These insights indicate strong generalization capabilities gained when applying our approach, which we argue is more relevant in practical scenarios of the AA task than in-domain. This also provides evidence of the importance of leveraging the various layers of the base transformer when working on the AA task. Furthermore, we perform experiments to analyze the role each layer plays in the overall performance of our model. We found that certain layers became more relevant when models are tested in the out-of-domain setting, leading to high performance. The code for our experiments will be publicly available upon acceptance.
@@ 2 Related Work
Author Attribution. Traditional methods rely on manually designed features, such as syntactic patterns and function word frequencies, to capture distinctive writing styles Koppel and Schler (2004). Several datasets were constructed to study the task Potthast et al. (2016). With the rise of deep learning, transformer-based models have set new records by leveraging their ability to capture nuanced stylistic patterns in text. For instance, Rivera-Soto et al. (2021) employed a contrastive learning approach to map texts written by the same author into similar regions of an embedding space. Similarly, Wegmann et al. (2022) introduced methods to ensure that embeddings focus on style rather than content, allowing for more reliable attribution. Despite their success, these approaches rely only on the final layer of the transformer model as the source of text representation. In contrast, our approach leverages different text representations from all the transformer’s layers. By applying our approach to these two aforementioned AA models, we demonstrate the gain from modeling the task at different linguistic granularities.
@@ 3 Approach
We propose a multi-layer contrastive learning approach that explicitly models authorship attribution at varying levels of linguistic granularity. Instead of relying on a single representation derived from the final layer, our approach extracts embeddings from all the N layers of the transformer, leveraging its linguistic capacity for authorship attribution.
As highlighted in Figure 1, our approach is transformer-based Vaswani et al. (2017) with additional N+1 Feed Forward layers that project each of the transformer’s hidden states into a new embedding space, including the initial token embeddings. We then apply a contrastive loss function on each of these N+1 embedding spaces to pull documents written by the same author together while keeping other documents further apart in the corresponding space. At each training step, the average loss of all the N contrastive losses is computed and back-propagated throughout the model. We hypothesize that this architecture ensures that stylistic features across multiple linguistic levels contribute to the final representation, allowing the model to learn complementary signals from different layers. Rather than collapsing stylistic information into a single embedding, we maintain separate representations for each layer. In Section 6.1, we present an analysis confirming how the various learned layers of our trained model capture different linguistic representations from texts. During inference, given two texts, we compute their similarity at each learned layer projection. These layer-wise similarity scores are aggregated to produce a final similarity measure that determines the likelihood of these two texts being written by the same author.
4 Experiment setup
@@
Table 1: Evaluation results comparing the effect of our approach (prefixed with LIGHT) when applied to two AA models, LUAR and Wegmann. Results are shown when all models are evaluated on the three datasets for in-domain (highlighted grey) and out-of-domain settings. The highest value for each model in each setting is highlighted in bold. The overall highest value among all models for each setting is highlighted with underlines.
In our evaluation, we closely follow the experiment setup of Rivera-Soto et al. (2021) in terms of task definition, datasets, and metrics used. To obtain reliable insights, we test our approach by applying it to two popular AA models: LUAR Rivera-Soto et al. (2021) and Wegmann Wegmann et al. (2022), considering two base transformers, RoBERTa Liu (2019) and ModernBERT Warner et al. (2024).
@@
We evaluate the effectiveness of our model on the Reddit, Amazon and Fanfiction datasets used in Rivera-Soto et al. (2021). These datasets represent distinct domains with different linguistic styles, levels of formality, and authorial consistency. Each dataset is partitioned into training and evaluation, ensuring that evaluation texts are temporally disjoint from training data where timestamps are available Andrews and Bishop (2019). For Amazon, we use 135K authors with at least 100 reviews, training on 100K, and evaluating on 35K Ni et al. (2019). The Fanfiction dataset Bevendorff et al. (2020) consists of 278,169 stories from 41K authors, with evaluation restricted to 16,456 authors, each contributing exactly two stories Bischoff et al. (2020). The Reddit dataset includes 1 million authors, with training data from Khan et al. (2021) and evaluation using a disjoint, temporally future set Baumgartner et al. (2020). Further information can be found in Appendix A.3.
@@
We follow previous work in evaluating all models using Recall-at-8 (R@8) and Mean Reciprocal Rank (MRR). Recall-at-8 measures the probability that the correct author appears among the top 8 ranked candidates, while MRR evaluates ranking quality based on the position of the first correct author match. Higher values indicate stronger performance Voorhees and Tice (2000).
AA Models.
@@
As mentioned, we apply our approach by extending two popular AA models, LUAR and Wegmann. LUAR Rivera-Soto et al. (2021) is a contrastive learning-based authorship attribution model with a RoBERTa model (paraphrase-distilroberta checkpoint [^1] https://huggingface.co/distilbert/distilroberta-base) as its base transformer. Given the two text collections, the model samples excerpts of 32 tokens and passes them through a RoBERTa transformer, obtaining only the final layer representation of each excerpt. An attention mechanism and max pooling layer are applied to aggregate the extracted representations into a single embedding representing each text collection. A contrastive loss is then applied to these representations to make them closer for positive pairs (written by the same author) and further apart for negative pairs. The Wegmann Wegmann et al. (2022) model uses a sentence transformer architecture trained with triplet loss over contrastive anchor-positive-negative sentence triplets to capture authorial style. The core idea of Wegmann is to construct challenging triplets where the negative sentence is from a different author but topically very similar to the anchor author to force the model to learn stylistic differences rather than topical ones. The model is trained on the Reddit dataset, making use of the Subreddit metadata as topical information. We modify each model by applying our approach to extend its architecture with N+1 projection layers. We also consider replacing the RoBERTa transformer base with ModernBERT, a more recent transformer architecture that allows a longer context window (up to 8k tokens) and is trained on a bigger dataset, leading to better performance on various tasks. This also allows us to make sure that our results apply to different transformer architectures. Additionally, we evaluate against the same baselines presented in Rivera-Soto et al. (2021). The first is a convolutional model Andrews and Bishop (2019), which encodes text representations using subword convolutional networks rather than transformers. The second baseline is a TF-IDF baseline, which represents documents as bag-of-words vectors weighted by term frequency in the text and normalized by inverse document frequency.
@@
Training Dataset Evaluation Dataset Model Reddit $\cup$ Fanfic Reddit $\cup$ Amazon Amazon $\cup$ Fanfic R@8 MRR R@8 MRR R@8 MRR Reddit LIGHT-LUAR 64.13 49.46 63.96 49.44 29.95 20.62 LUAR 56.35 41.20 60.58 45.46 20.40 13.11 Amazon LIGHT-LUAR 69.43 57.04 89.00 78.38 85.99 74.23 LUAR 53.51 39.63 84.84 72.09 79.60 64.83 Fanfic LIGHT-LUAR 54.16 44.56 43.24 33.54 51.87 42.63 LUAR 57.51 46.32 40.69 30.49 51.84 41.79
@@
Table 2: Evaluation Results after training on multiple domain datasets shown only for LIGHT-LUAR and LUAR
Training.
For LUAR, we follow the training process of Rivera-Soto et al. (2021) to train a reproduced version of their LUAR model with RoBERTa base. We train a model for each of the three datasets for 20 epochs, with validation performed every five epochs. We use a batch size of 84 and sample from each author 16 text excerpts of 32 tokens each. For the ModernBERT-based version, we extend training to 30 epochs and 32 text excerpts to accommodate the model’s larger size and longer context capacity, respectively, while keeping the excerpt length fixed at 32 tokens and maintaining validation every five epochs. As for Wegmann, we fine-tune the sentence transformer model using triplet loss over 58k triplets constructed from the Reddit training dataset following Wegmann’s sampling procedure. Since such topical information is not provided in Amazon or Fanfiction datasets, we refrain from training on them. We train separate models with either RoBERTa or ModernBERT as base transformer, using a batch size of 24 for 20 epochs. Our proposed approach (prefixed as LIGHT) extends this by retrieving hidden states from all transformer layers and computing triplet-loss at each layer individually, averaging the resulting losses to promote consistent style alignment throughout the network.
5 Results
We examine the impact of our approach when training on single and multiple domains. In the following, we focus on presenting our results considering RoBERTa as the base transformer. Results when using ModernBERT are presented in Appendix B due to space limitations. Note that the findings presented here also apply to ModernBERT.
@@ 5 Results
Figure 2: The linguistic capability of each of the 6 projected layers of the LIGHT-LUAR model trained on the Reddit dataset compared to the output layer of the LUAR model (first column) as probed using Holmes benchmark.
5.1 Single-Domain Training
In the single-domain setting, models are trained on a single dataset and tested on the test split of all three datasets to assess their generalizability. Table 1 shows the recall at 8 (R@8) and mean reciprocal rank (MRR) scores. Since training Wegmann’s model requires topical information that only exists in the Reddit dataset, the model and our extended version of it (LIGHT-Wegmann) are only trained on Reddit and evaluated on all three datasets.
In general, LUAR and our LIGHT-LUAR model outperform all other baselines in all scenarios except when models are trained on Reddit and evaluated on Amazon. For in-domain evaluation, we observe that applying our approach to LUAR and Wegmann leads to better results only on the Amazon dataset for LUAR and Reddit for Wegmann. In all other in-domain scenarios, applying our approach leads to slightly worse performance. This suggests that for in-domain evaluation, final-layer representations contain sufficient information for authorship attribution, and incorporating multiple layers might lead to a slight degradation in the performance. We argue that the improvement on the Amazon dataset is due to the structured writing style, allowing the different layers to capture more nuanced linguistic representations from texts. In contrast, informal and discussion-based datasets, such as Reddit, exhibit less improvement, likely due to the topic-driven nature of the writing, which can overshadow stylistic signals. This is not the case for LIGHT-Wegmann, which outperforms its baseline (Wegmann) since the model is trained to better separate content from style.
For out-of-domain evaluation, applying our approach into both AA models demonstrates substantial improvements over baselines, significantly enhancing cross-domain generalization and establishing new state-of-the-art results. LIGHT-LUAR trained on homogeneous datasets, such as Amazon, show remarkable adaptability to free-form writing (An MRR of 18.03 for LIGHT-LUAR compared to 14.65 for LUAR on Reddit dataset), while those trained on Reddit generalize exceptionally well to both Amazon and Fanfiction. However, applying our approach to the Wegmann model leads to improvement in performance in the case of Fanfic but not Amazon. In case of LIGHT-Wegmann with ModernBERT (Table 3 in the appendix), our approach is better in all out-of-domain cases. Compared to original AA models, our approach substantially improves transferability, highlighting the importance of leveraging different linguistic granularity levels rather than relying solely on final-layer representations. Most notably, in the Fanfiction transfer setting, where LUAR struggles, our approach achieves 50% R@8 improvements.
5.2 Multi-domain Training